OpenStudy (anonymous):

A building lot in a city is shaped as a 30° -60° -90° triangle. The side opposite the 30° angle measures 41 feet. a. Find the length of the side of the lot opposite the 60° angle. b. Find the length of the hypotenuse of the triangular lot. c. Find the sine, cosine, and tangent of the 30° angle in the lot. Write your answers as

OpenStudy (anonymous):



OpenStudy (anonymous):

A---->\[41\sqrt{3}\]

OpenStudy (anonymous):

lenth of hypotenuse B------>82

OpenStudy (anonymous):

sin30=1/2, cos30= (root 3)/2 and tan 30=1/(root 3)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

my pleasure!!